OpenStudy (anonymous):

Simplify: \[(3z)^{2} (6z^{2})^{-3} \]

OpenStudy (anonymous):

\[(3z ^{2})(6z ^{2})^{-3}=(3z ^{2})6^{-3} z ^{-6}=3*6^{-3}*z ^{-4}=z ^{-4}/108\]

OpenStudy (anonymous):

answer is \[{1 \over 24z^{4}}\]

OpenStudy (anonymous):

sorry your prob was (3z)^2...and i did it with 3z^2....so i got wrong...u r write..

OpenStudy (anonymous):

can you explain how to get the answer?

OpenStudy (anonymous):

\[(3z)^2(6z^2)^{-3}=(9z^2)/(6z^2)^3=\frac{9z^2}{216z^6}=1/24z^4\]

OpenStudy (anonymous):

okay. okay... i'm still trying to grasp this

OpenStudy (anonymous):

wait... it's not making sense. why divide on the second step?

OpenStudy (anonymous):

where did you get 216z^6 on the third step?

OpenStudy (anonymous):

a^-x=1/(a^x) do u know that rule?

OpenStudy (anonymous):

nope. haa

OpenStudy (anonymous):

(6z^2)^3=6^3z^6=216z^6

OpenStudy (anonymous):

okay i see where the divde comes from. the negative exponent.